# Instruction Tuning With Loss Over Instructions

### v1

@@ Abstract
Instruction tuning plays a crucial role in shaping the outputs of language models (LMs) to desired styles. In this work, we propose a simple yet effective method, Instruction Modelling (IM), which trains LMs by applying a loss function to the instruction and prompt part rather than solely to the output part. Through experiments across 21 diverse benchmarks, we show that, in many scenarios, IM can effectively improve the LM performance on both NLP tasks (e.g., MMLU, TruthfulQA, and HumanEval) and open-ended generation benchmarks (e.g., MT-Bench and AlpacaEval). Remarkably, in the most advantageous case, IM boosts model performance on AlpacaEval 1.0 by over 100%. We identify two key factors influencing the effectiveness of IM: (1) The ratio between instruction length and output length in the training data; and (2) The number of training examples. We observe that IM is especially beneficial when trained on datasets with lengthy instructions paired with brief outputs, or under the Superficial Alignment Hypothesis (SAH) where a small amount of training examples are used for instruction tuning. Further analysis substantiates our hypothesis that our improvement can be attributed to reduced overfitting to instruction tuning datasets. Our work provides practical guidance for instruction tuning LMs, especially in low-resource scenarios. Our code is available at https://github.com/ZhengxiangShi/InstructionModelling.
@@ 1 Introduction
Figure 1: Performance differences between Instruction Tuning (IT) and our proposed method instruction modelling (IM) trained on 7 instruction tuning datasets. (Left) The mean performance across 18 traditional NLP tasks. (Right) The win rate on the AlpacaEval 1.0 benchmark. Please refer to §4.2 for details.
Language models (LMs) are trained to predict the next token on massive corpora, enabling them to learn general-purpose representations transferable to various language understanding or generation tasks [7, 47, 48, 56, 61]. However, it does align LMs to act in accordance with the user’s intentions [34]. To enable this transfer, various methods for aligning language models [3, 14, 44, 72, 78] have thus been proposed, one of which is instruction tuning (IT) [36, 64, 68]. Recent study [77] proposes Superficial Alignment Hypothesis (SAH): A model’s knowledge and capabilities are learnt almost entirely during pretraining, only minimal instruction tuning data is required to enable high-quality outputs in the desired output style. Existing works [1, 43, 44, 51, 64, 68, 36] mainly perform instruction tuning by focusing the loss computation solely on the output segments.
Figure 2: (Left) Performance improvement, achieved by our approach instruction modelling (IM) compared to Instruction Tuning (IT) on the AlpacaEval 1.0, against the ratio between average instruction length and average output length in instruction tuning datasets (training size noted in parentheses). We highlight several representative instruction tuning datasets in yellow. Our analysis suggests that IM is especially beneficial for datasets characterized by lengthy instructions or prompts paired with comparably brief outputs, such as Code Alpaca [10] and and Less MMLU Chat [67]. (Right) Performance improvement achieved by our approach IM over IT on the AlpacaEval 1.0 against the number of training examples in instruction tuning datasets. Here we maintain a fixed ratio between instruction and output length of 10. This analysis suggests that IM is particularly effective under the low-resource setting or Superficial Alignment Hypothesis. Please refer to §4.2 for details.
@@
  • We propose Instruction Modelling (IM), aiming to enhance both the instruction-following and general performance on NLP tasks of LMs. Through extensive experiments across 21 benchmarks, we demonstrate that, in many scenarios, IM substantially improves performance of LMs trained on various instruction tuning datasets, particularly notable in the AlpacaEval 1.0 benchmark where it boosts scores by over 100%.
  • Our study identifies key factors influencing the effectiveness of IM, including the ratio between instruction length and output length and the number of training examples, providing practical guidance for instruction tuning LMs, especially under the low-resource scenarios.
@@ Conclusion.
In conclusion, our study proposes Instruction Modelling, which trains LMs with loss over instructions rather than outputs only. Our experimental evaluations demonstrate that our approach largely improves the performance of LMs on both NLP tasks and open-ended generation benchmarks in some scenarios, especially under the Superficial Alignment Hypothesis and low-resource setting where minimal training data is used for instruction tuning. Our analysis has shed light on two key factors that influence the effectiveness of our approach, (1) the ratio between instruction and output lengths, and (2) the quantity of training data, providing practical insights for optimising instruction-based training methods. Additionally, our analysis reveals the mechanisms behind the effectiveness of IM, particularly its ability to reduce overfitting, showing that applying instruction losses in some scenarios can lead to more robust and adaptable LMs.
@@
Here we discuss some potential limitations and the broader impact of our work. Several limitations are outlined as follows: (1) The success of our approach relies on the quality and diversity of the instructions and prompts in the training datasets. Poorly defined or ambiguous instructions may undermine the effectiveness of IM, leading to sub-optimal performance. (2) It is crucial to ensure that the instructions are ethically sound and free from harmful or biased content. Training on inappropriate or toxic instructions may result in undesirable outputs. Previous works [5, 7, 4] have extensively discussed the risks and potential harms associated with LMs, including the amplification of undesirable biases learned from training data [5, 2, 9]. Our work has the potential to positively impact the community by helping to mitigate overfitting, resulting in models that are more robust and generalise better to new data, especially in low-resource scenarios. This can enhance the reliability and trustworthiness of AI systems in real-world applications.
@@
provides experimental setting, implementation details and hyperparameters for all comparison methods used in our experiments.
@@ Appendix B Evaluation Datasets and Details
The model evaluation utilizes the dataset at the huggingface dataset  [^7] https://huggingface.co/datasets/lukaemon/bbh, specifically tested on the ‘test‘ split without the use of few-shot examples. We follow the setup in previous works [30, 59]. The evaluation metric is the exact match score, averaged (mean) to assess performance. Generation is constrained to a maximum of 1024 tokens, with termination upon encountering specific delimiters such as "</s>", "Q", or double newlines. The generation is greedy decoding (temperature set to 0.0) and does not use sampling. Answer extraction employs regex patterns to identify responses immediately following "the answer is" and captures only the first occurrence.
@@
The evaluation setup for the dataset at the huggingface dataset  [^9] https://huggingface.co/datasets/allenai/ai2_arc utilizes a multiple-choice format. We follow the HuggingFace Open LLM Leaderboard to 25 few-shot examples. The performance metric used is mean normalized accuracy (acc_norm).
@@ Appendix C Implementation Details
Here we provide a detailed experimental design for Figure 2 Right. We strategically sampled varying sizes of training examples from the Tulu V2 dataset to investigate the effectiveness of IM with different sizes training examples. Starting with approximately 320,000 examples in the Tulu V2 dataset, we creates subsets of data ranging from as few as 1,000 to as many as 35,000 examples. These subsets were selected randomly, ensuring a representative mix across different scales. We adhered to a fixed instruction-to-output length ratio of approximately 10 to maintain consistency in training conditions across all samples. We train the Llama-2-7B-Base on all these subsets and evaluate them respectively.
@@ Appendix E The impact of Epochs on the Win Rate
The figure 10 illustrates the comparative analysis of AlpacaEval 1.0 scores across different epochs for two datasets, LIMA and Alpagasus Dolly 9k datasets. We evaluate the performance of IM and IT over different numbers of epochs. IM consistently surpasses IT in performance on the Alpagasus Dolly 9k dataset, while the performance of both approaches is comparable on the LIMA dataset.
@@ Appendix F Applying KL Divergence Loss for Instruction Tuning
Kullback-Leibler (KL) divergence is commonly employed as a regularisation method in the fine-tuning of LMs, helping to mitigate overfitting by constraining the fine-tuned model to remain similar to the pre-trained model [44]. Specifically, the KL divergence is added to the fine-tuning objective as a per-token regularisation term between the fine-tuned LM $\pi_{\theta}(x)$, and the pre-trained LM, $\pi^{\text{pre}}(x)$. For supervised fine-tuning with next token prediction loss, the training objective incorporating KL divergence is computed as follows:
@@ Appendix G The impact of Epochs on Output Lengths
Figure 11 illustrate the average output length of various models across different epochs. We report the output length on four different datasets, including Alpagasus Dolly 3k, Alpagasus Dolly 9k, LIMA, and Less Tydiqa. Each line represents the average output length of a model, with epochs ranging from 2 to 10, and is accompanied by error bars that denote the normalised standard deviation (10%) of the output lengths. Our experimental results show that our approach IM does not consistently increase the output length and that win rates are not necessarily associated with the length of the output.

### v2

@@ Abstract
Instruction tuning plays a crucial role in shaping the outputs of language models (LMs) to desired styles. In this work, we propose a simple yet effective method, Instruction Modelling (IM), which trains LMs by applying a loss function to the instruction and prompt part rather than solely to the output part. Through experiments across 21 diverse benchmarks, we show that, in many scenarios, IM can effectively improve the LM performance on both NLP tasks (e.g., MMLU, TruthfulQA, and HumanEval) and open-ended generation benchmarks (e.g., MT-Bench and AlpacaEval). Remarkably, in the most advantageous case, IM boosts model performance on AlpacaEval 1.0 by over 100%. We identify two key factors influencing the effectiveness of IM: (1) The ratio between instruction length and output length in the training data; and (2) The number of training examples. We observe that IM is especially beneficial when trained on datasets with lengthy instructions paired with brief outputs, or under the Superficial Alignment Hypothesis (SAH) where a small amount of training examples are used for instruction tuning. Further analysis substantiates our hypothesis that our improvement can be attributed to reduced overfitting to instruction tuning datasets. It is worth noting that we are not proposing IM as a replacement for current fine-tuning processes. Instead, our work aims to provide practical guidance for instruction tuning LMs, especially in low-resource scenarios. Our code is available at https://github.com/ZhengxiangShi/InstructionModelling.
@@ 1 Introduction
Figure 1: Performance differences between Instruction Tuning (IT) and our proposed method instruction modelling (IM) trained on 7 instruction tuning datasets. These datasets contain prompts and responses but do not contain preference pairs. Specifically, we use the Less datasets [68] and Alpagasus datasets [11], which are subsets of Flan V2 [14], Dolly [18], and Stanford Alpaca [61] to ensure good performance. We also report the results on the LIMA dataset. (Left) The mean performance across 18 traditional NLP tasks (see §4.1 for details). (Right) The win rate on the AlpacaEval 1.0 benchmark [37]. Please refer to §4.2 for details.
Language models (LMs) are trained to predict the next token on massive corpora, enabling them to learn general-purpose representations transferable to various language understanding or generation tasks [7, 47, 48, 56, 62]. However, it does align LMs to act in accordance with the user’s intentions [34]. To enable this transfer, various methods for aligning language models [3, 14, 44, 73, 58, 79] have thus been proposed, one of which is instruction tuning (IT) [36, 65, 69]. Recent study [78] proposes Superficial Alignment Hypothesis (SAH): A model’s knowledge and capabilities are learnt almost entirely during pretraining, only minimal instruction tuning data is required to enable high-quality outputs in the desired output style. Existing works [1, 43, 44, 51, 65, 69, 36] mainly perform instruction tuning by focusing the loss computation solely on the output segments.
Figure 2: (Left) Performance improvement, achieved by our approach instruction modelling (IM) compared to Instruction Tuning (IT) on the AlpacaEval 1.0, against the ratio between average instruction length and average output length in instruction tuning datasets (training size noted in parentheses). We highlight several representative instruction tuning datasets in yellow. Our analysis suggests that IM is especially beneficial for datasets characterized by lengthy instructions or prompts paired with comparably brief outputs, such as Code Alpaca [10] and Less MMLU Chat [68]. (Right) Performance improvement achieved by our approach IM over IT on the AlpacaEval 1.0 against the number of training examples in instruction tuning datasets. Here we maintain a fixed ratio between instruction and output length of 10. This analysis suggests that IM is particularly effective under the low-resource setting or Superficial Alignment Hypothesis. Please refer to §4.2 for details.
@@
  • We propose Instruction Modelling (IM), aiming to enhance both the instruction-following and general performance on NLP tasks of LMs. Through extensive experiments across 21 benchmarks, we demonstrate that, in many scenarios, IM substantially improves the performance of LMs trained on various instruction tuning datasets, particularly notable in the AlpacaEval 1.0 benchmark where it boosts scores by over 100%.
  • Our study identifies key factors influencing the effectiveness of IM, including the ratio between instruction length and output length and the number of training examples. We are not proposing IM as a replacement for current instruction tuning processes. Rather, we provide empirical guidance for fine-tuning LMs, especially under low-resource scenarios.
@@ Conclusion.
Our study proposes Instruction Modelling, which trains LMs with loss over instructions rather than outputs only. Our experimental evaluations demonstrate that our approach largely improves the performance of LMs on both NLP tasks and open-ended generation benchmarks in some scenarios, especially under the Superficial Alignment Hypothesis and low-resource setting where minimal training data is used for instruction tuning. Our analysis has shed light on two key factors that influence the effectiveness of our approach, (1) the ratio between instruction and output lengths, and (2) the quantity of training data, providing practical insights for optimising instruction-based training methods. Additionally, our analysis reveals the mechanisms behind the effectiveness of IM, particularly its ability to reduce overfitting, showing that applying instruction losses in some scenarios can lead to more robust and adaptable LMs.
@@
Here we discuss some potential limitations and the broader impact of our work. Several limitations are outlined as follows: (1) The success of our approach relies on the quality and diversity of the instructions and prompts in the training datasets. Poorly defined or ambiguous instructions may undermine the effectiveness of IM, leading to sub-optimal performance; and (2) It is crucial to ensure that the instructions are ethically sound and free from harmful or biased content. Training on inappropriate or toxic instructions may result in undesirable outputs. Previous works [5, 7, 4] have extensively discussed the risks and potential harms associated with LMs, including the amplification of undesirable biases learned from training data [5, 2, 9]. Our work has the potential to positively impact the community by helping to mitigate overfitting, resulting in models that are more robust and generalise better to new data, especially in low-resource scenarios. This can enhance the reliability and trustworthiness of AI systems in real-world applications.
Acknowledgments and Disclosure of Funding
The authors express their gratitude to the NeurIPS reviewers and area chairs for their insightful discussions. Zhengyan is funded by the Research Studentship from University College London (UCL). Bin is funded by the Bloomberg Fellowship. The authors gratefully appreciate the generous support from OpenAI for providing API credits through the OpenAI API Researcher Access Program.
@@
provides the experimental setting, implementation details and hyperparameters for all comparison methods used in our experiments.
@@ Appendix B Evaluation Datasets and Details
The model evaluation utilises the dataset at the huggingface dataset  [^7] https://huggingface.co/datasets/lukaemon/bbh, specifically tested on the ‘test‘ split without the use of few-shot examples. We follow the setup in previous works [30, 60]. The evaluation metric is the exact match score, averaged (mean) to assess performance. Generation is constrained to a maximum of 1024 tokens, with termination upon encountering specific delimiters such as "</s>", "Q", or double newlines. The generation is greedy decoding (temperature set to 0.0) and does not use sampling. Answer extraction employs regex patterns to identify responses immediately following "the answer is" and captures only the first occurrence.
@@
The evaluation setup for the dataset at the huggingface dataset  [^9] https://huggingface.co/datasets/allenai/ai2_arc utilises a multiple-choice format. We follow the HuggingFace Open LLM Leaderboard to 25 few-shot examples. The performance metric used is mean normalised accuracy (acc_norm).
@@ Appendix C Implementation Details
Here we provide a detailed experimental design for Figure 2 Right. We strategically sampled varying sizes of training examples from the Tulu V2 dataset to investigate the effectiveness of IM with different sizes of training examples. Starting with approximately 320,000 examples in the Tulu V2 dataset, we create subsets of data ranging from as few as 1,000 to as many as 35,000 examples. These subsets were selected randomly, ensuring a representative mix across different scales. We adhered to a fixed instruction-to-output length ratio of approximately 10 to maintain consistency in training conditions across all samples. We train the Llama-2-7B-Base on all these subsets and evaluate them respectively.
@@ Appendix E The impact of Epochs on the Win Rate
Figure 10 illustrates the comparative analysis of AlpacaEval 1.0 scores across different epochs for two datasets, LIMA and Alpagasus Dolly 9k datasets. We evaluate the performance of IM and IT over different numbers of epochs. IM consistently surpasses IT in performance on the Alpagasus Dolly 9k dataset, while the performance of both approaches is comparable on the LIMA dataset.
@@ Appendix F Applying KL Divergence Loss for Instruction Tuning
Kullback-Leibler (KL) divergence is commonly employed as a regularisation method in the fine-tuning of LMs, helping to mitigate overfitting by constraining the fine-tuned model to remain similar to the pre-trained model [44]. Specifically, the KL divergence is added to the fine-tuning objective as a per-token regularisation term between the fine-tuned LM $\pi_{\theta}(x)$, and the pre-trained LM, $\pi^{\text{pre}}(x)$. For supervised fine-tuning with the next token prediction loss, the training objective incorporating KL divergence is computed as follows:
@@ Appendix G The impact of Epochs on Output Lengths
Figure 11 illustrates the average output length of various models across different epochs. We report the output length on four different datasets, including Alpagasus Dolly 3k, Alpagasus Dolly 9k, LIMA, and Less Tydiqa. Each line represents the average output length of a model, with epochs ranging from 2 to 10, and is accompanied by error bars that denote the normalised standard deviation (10%) of the output lengths. Our experimental results show that our approach IM does not consistently increase the output length and that win rates are not necessarily associated with the length of the output.